OpenStudy (anonymous):

Solve 7x - 4 = 5x + 15 for x.

OpenStudy (midhun.madhu1987):

Move all terms containing x to the left side and numbers to right side and simplify.

OpenStudy (anonymous):

7x-5x=15+4 2x=19 x=19/2=9.5

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

7(x - 2) = 3(x + 4)